2x cubed + 2= 0
I dont understand how 2x breaks down.

Answer by stanbon(48558)

2x cubed + 2= 0
2x^3 = -2
x^3 = -1
x^3+1 = 0
(x+1)(x^2-x+1) = 0
x = -1 or x = [1 +- sqrt(1-4)]/2
x = -1 or x = [1+isqrt(3)]/2 or x = [1-isqrt(3)]/2
============
Cheers,
Stan H.